OpenStudy (anonymous):

Which fraction falls between 3/10 and 3/5 A.5/10 B.8/10 C.4/5 D.2/10

OpenStudy (anonymous):

I got C is that even near right?

OpenStudy (anonymous):

So you can convert all the fractions into decimals and see which option lies between 3 / 5 and 3 / 10

OpenStudy (anonymous):

wait how do i convert them to decimals?

OpenStudy (anonymous):

C is wrong by the way

OpenStudy (anonymous):

yah i know

OpenStudy (anonymous):

Make the denominator 10. Example 1 / 10 = 0.1 2 / 10 = 0.2 3 / 10 = 0.3 etc

OpenStudy (anonymous):

oh ok im going to try and see if i get it

OpenStudy (anonymous):

i got A... is that right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thanks